OpenStudy (anonymous):

Which of these is the algebraic expression for "five more than the product of ten and some number?" 10 + 5x 5 + 10 + x 10x + 5 5 + 10 ÷ x

OpenStudy (anonymous):

_x is times it

OpenStudy (ltrout):

5 + 10x

OpenStudy (anonymous):

b

OpenStudy (anonymous):

?

OpenStudy (ltrout):

No. That is the sum of 5 plus 10 plus a number. You want 10 times a number so 10x.

OpenStudy (anonymous):

but if it was x wouldnt it say 5 tmes not mre than?

OpenStudy (ltrout):

product of ten and a number = 10 * x or 10x. Sum of 5 and the product of ten and a number = 5 + 10 * x = 5 + 10x || 10x + 5.

OpenStudy (anonymous):

ok so ill get c

OpenStudy (ltrout):

Yep :)

OpenStudy (anonymous):

ok hmm thxxs

OpenStudy (ltrout):

Care to medal?

OpenStudy (anonymous):

idk how im new

OpenStudy (ltrout):

See that blue button next to my responses that says "Best Response"? Click that

OpenStudy (anonymous):

ok but i need to see if its right u know wht the flutter ill give it to u

OpenStudy (ltrout):

I can promise you it's right. And if you ever want to take it away after the fact, there is now an "x" that says "Undo"

OpenStudy (anonymous):

ok